OpenStudy (anonymous):

Completely factor the expression. x2 + 12x + 27

OpenStudy (wolf1728):

The 27 should give you a clue. (9 x 3 = 27)

OpenStudy (anonymous):

@Qdogmayne:Do u want geting factor from this ?:)

OpenStudy (anonymous):

ok but what about x^2 +12x

OpenStudy (anonymous):

split the mid term

OpenStudy (anonymous):

In get factor allllllllways u should get the same things for example x at there ! OK ?!

OpenStudy (anonymous):

or else take the common mulyitple of all numbers out and factorise them

OpenStudy (wolf1728):

As I said earlier the 27 should give you a clue 9 x 3 = 27 AND 9 + 3 = 12

OpenStudy (anonymous):

Now look : first one wxample : 2.8+2.9=2(8+9) Now : X^2/x+12x/x=x(x+12) Got it ?!:)

OpenStudy (anonymous):

@Qdogmayne

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

So whats the problem ? Any question ?!

OpenStudy (wolf1728):

Odogmayne - could you post your answer ? (just to make sure it is right)

OpenStudy (anonymous):

(x+3)(x+9

OpenStudy (wolf1728):

We were waiting for Odogmayne to type something in. Now everyone's gone.

OpenStudy (anonymous):

ok so I'm doin 2(8+9)?